# UAV-UGV Cooperative Trajectory Optimization and Task Allocation for Medical Rescue Tasks in Post-Disaster Environments

### v1

@@ Abstract
In post-disaster scenarios, rapid and efficient delivery of medical resources is critical and challenging due to severe damage to infrastructure. To provide an optimized solution, we propose a cooperative trajectory optimization and task allocation framework leveraging unmanned aerial vehicles (UAVs) and unmanned ground vehicles (UGVs). This study integrates a Genetic Algorithm (GA) for efficient task allocation among multiple UAVs and UGVs, and employs an informed-RRT* (Rapidly-exploring Random Tree Star) algorithm for collision-free trajectory generation. Further optimization of task sequencing and path efficiency is conducted using Covariance Matrix Adaptation Evolution Strategy (CMA-ES). Simulation experiments conducted in a realistic post-disaster environment demonstrate that our proposed approach significantly improves the overall efficiency of medical rescue operations compared to traditional strategies, showing substantial reductions in total mission completion time and traveled distance. Additionally, the cooperative utilization of UAVs and UGVs effectively balances their complementary advantages, highlighting the system’s scalability and practicality for real-world deployment.
Index Terms:
@@
I Introduction
Earthquakes, floods, hurricanes, and other catastrophic events often destroy roads, bridges, and communication networks, severely limiting traditional emergency response capabilities. Consequently, ensuring timely medical assistance becomes an urgent priority, directly influencing survival rates and recovery outcomes. Gür et al.[1] assess the emergency response to the 2023 Kahramanmaras earthquakes (7.7 and 7.6 magnitude), noting that severe infrastructure damage impeded rescue efficiency and medical aid, adversely affecting survival rates and recovery, emphasizing the need for better preparedness and resilient infrastructure.
UAVs and UGVs have emerged as promising solutions due to their flexibility, rapid deployment, and ability to navigate environments that are hazardous or inaccessible to human responders. UAVs offer advantages such as aerial reconnaissance, swift delivery, and the ability to bypass ground-level obstacles, while UGVs provide stable transportation of heavier supplies and longer operational endurance.
However, effective cooperation between UsAVs and UGVs presents several critical challenges. These include optimal task allocation among multiple vehicles with varying capabilities and constraints, robust and real-time trajectory planning in uncertain environments filled with obstacles, and the need to maintain efficient coordination to minimize response time and maximize resource utilization. Addressing these challenges is essential to enhance rescue efficiency and ultimately save lives.
@@ I Introduction
Multi-robot cooperative systems involving UAVs and UGVs have received significant attention as a promising approach for disaster relief operations. Cooperation between these platforms leverages their complementary strengths, allowing UAVs to quickly survey inaccessible regions and rapidly deliver critical medical supplies, while UGVs provide consistent and sustained logistical support on the ground. For instance, de Castro et al. [2] investigated UAV-UGV cooperation for coverage tasks in dynamic environments, while Krizmancic et al. [3] focused on coordination for construction tasks, and Cognetti et al. [4] utilized UAVs for localization to aid UGV formation and navigation. The relevance of such systems in Synthetic Aperture Radar operations has been reviewed by Queralta et al. [5].
@@
Effective task allocation methods are crucial in disaster scenarios to ensure optimal resource deployment. Techniques such as GA, Particle Swarm Optimization (PSO), and auction-based mechanisms have been studied extensively for their ability to efficiently allocate tasks under dynamic and uncertain conditions. For example, Li et al. [6] developed an IGA for multi-robot task assignment, and López-González et al. [7] applied parallel GAs for formation control. Dadgar et al. [8] proposed an adaptive PSO (A-RPSO) for target searching, while Elango et al. [9] combined K-means clustering with auction mechanisms for balanced task allocation. These methods enable multiple robots to dynamically adjust to changing environments, improving operational effectiveness.
@@
Path planning and optimization techniques also play a key role in enhancing the efficiency and safety of multi-robot rescue missions. Algorithms like RRT*, Probabilistic Roadmap (PRM), and A* search have been widely utilized to generate collision-free trajectories in complex, cluttered environments. Li et al. [10] specifically extended RRT* for multi-robot scenarios considering motion constraints. Other approaches like the PRM method have been combined with optimization techniques, as seen in the hybrid ABC-PRM-EP framework by Kumar and Sikander [11]. Additionally, trajectory optimization methods such as CMA-ES have been introduced to refine initial paths, minimizing travel distance and overall mission completion time. Díez and Aznar [12] compared CMA-ES with other evolution strategies for learning autonomous navigation behaviors in robot swarms. Integrating these task allocation and path planning strategies significantly contributes to the success of cooperative multi-robot disaster relief operations.
Although existing studies have made notable advancements in multi-robot cooperative systems, task allocation strategies, and trajectory optimization techniques for post-disaster rescue, several critical limitations remain unresolved. First, most current cooperative frameworks consider homogeneous teams of unmanned vehicles (either UAVs or UGVs alone), ignoring the potential benefits arising from heterogeneous cooperation. UAVs and UGVs each offer distinct capabilities—UAVs excel in agility and rapid coverage, while UGVs provide enhanced payload capacities and extended operational durations. However, these complementary strengths remain insufficiently integrated and exploited in current frameworks.
@@
Second, many task allocation methodologies previously proposed assume static environments and predetermined task sets. This assumption fails to reflect real-world disaster scenarios characterized by unpredictable task emergence, environmental dynamics, and complex obstacle distributions. Traditional optimization approaches and heuristic methods often struggle with the high computational complexity arising in such real-time, dynamic, multi-task scenarios.
@@
Third, trajectory optimization in disaster environments frequently focuses solely on feasibility or path length minimization without adequate consideration of smoothness, continuity, or energy efficiency constraints critical to actual UAV and UGV operations. Moreover, existing path planning algorithms typically produce jagged or inefficient trajectories, which are impractical for real-world navigation tasks.
@@
  • We propose an enhanced GA-based task allocation method that dynamically assigns rescue tasks among heterogeneous fleets of UAVs and UGVs, significantly improving task execution efficiency and operational synergy.
  • We integrate Informed-RRT* path planning with CMA-ES optimization to simultaneously achieve collision-free, efficient, and dynamically optimized trajectories in cluttered environments.
  • A systematic trajectory post-processing approach is applied to ensure practical smoothness and continuity, effectively addressing the shortcomings of previous planning frameworks and enhancing operational feasibility.
  • Through extensive simulation experiments based on realistic post-disaster scenarios, we validate the effectiveness, robustness, and scalability of our proposed approach, demonstrating significant improvements compared to existing methods.
II Problem Formulation
@@
II-A Environment Modeling
The operational theatre is conceptualized as a bounded two-dimensional Euclidean space, denoted by $\mathcal{E}\subset\mathbb{R}^{2}$, which represents a disaster-impacted zone, typically an urban or semi-urban landscape significantly damaged by events such as earthquakes or widespread flooding. Within this environment $\mathcal{E}$, several key elements are defined. A primary consideration is the presence of $N$ static obstacles, collected in the set $\mathcal{O}=\{O_{1},O_{2},\dots,O_{N}\}$. These obstacles, geometrically represented (e.g., as polygons or circles), correspond to impassable areas like collapsed structures, extensive debris fields, or flooded regions that impede vehicle movement. The locations and geometries of these major obstacles are assumed to be known or rapidly mappable post-disaster. Distributed across the environment are $M$ discrete task locations, $\mathcal{T}=\{T_{1},T_{2},\dots,T_{M}\}$, each signifying a point where a specific medical rescue action is required, such as the delivery of critical supplies or site assessment. For the scope of this work, these task locations are assumed to remain fixed throughout the planning phase. Finally, a central base station $B$ is situated within $\mathcal{E}$, serving as the designated origin and final destination for all unmanned vehicles involved in the operation, functioning as the mission control and logistics hub. Consequently, the navigable area for the vehicles, or free space, is $\mathcal{E}_{free}=\mathcal{E}\setminus\bigcup_{i=1}^{N}O_{i}$. A fundamental requirement is that all vehicle movement must be confined to $\mathcal{E}_{free}$, ensuring mandatory collision avoidance with respect to all obstacles in $O_{i}\in\mathcal{O}$. While real-world disaster zones are inherently dynamic, this study adopts a static representation of major obstacles and task points for planning tractability during a given mission cycle.
II-B UAV and UGV Operational Constraints
@@
where Pi are waypoints, $d(,)$ denotes the Euclidean distance. Kinematic limitations regarding velocity, turning radius, and acceleration also exist, though often addressed through simplified models or post-processing smoothing. Collision avoidance with environmental obstacles and other vehicles remains a critical constraint.
@@
II-C Task Definition and Constraints
@@ II-D Optimization Objectives
II-E Problem Scope and Formal Summary
In essence, the problem tackled in this paper is the optimization of task allocation and trajectory planning for a heterogeneous fleet $\mathcal{V}$ of UAVs and UGVs operating in a post-disaster setting. The operational context is defined by an environment $\mathcal{E}$ containing a set of known, static obstacles $\mathcal{O}$, a designated set of medical rescue task locations $\mathcal{T}$, and a central base station $B$. The core objective is to determine an optimal operational plan, which encompasses both the assignment of each task $T_{j}\in\mathcal{T}$ exclusively to a single vehicle $V_{i}\in\mathcal{V}$ (represented by assignment $X$), and the subsequent generation of an ordered task sequence $\pi_{i}$ and a feasible, collision-free trajectory $\mathcal{P}_{i}(t)$ for each vehicle. Each trajectory must originate from $B$, visit the assigned tasks in the determined sequence $\pi_{i}$, and terminate back at $B$, while strictly adhering to all vehicle-specific constraints, particularly maximum travel endurance and payload limits. The quality of the plan is primarily evaluated based on minimizing the overall mission completion time and the total distance traveled by the entire fleet.

### v2

@@ Abstract
In post-disaster scenarios, rapid and efficient delivery of medical resources is critical and challenging due to severe damage to infrastructure. To provide an optimized solution, we propose a cooperative trajectory optimization and task allocation framework leveraging unmanned aerial vehicles (UAVs) and unmanned ground vehicles (UGVs). This study integrates a Genetic Algorithm (GA) for efficient task allocation among multiple UAVs and UGVs, and employs an informed-RRT* (Rapidly-exploring Random Tree Star) algorithm for collision-free trajectory generation. Further optimization of task sequencing and path efficiency is conducted using Covariance Matrix Adaptation Evolution Strategy (CMA-ES). Simulation experiments conducted in a realistic post-disaster environment demonstrate that our proposed approach significantly improves the overall efficiency of medical rescue operations compared to traditional strategies. Specifically, our method reduces the total mission completion time to 26.7 minutes for a 15-task scenario, outperforming K-Means clustering and random allocation by over 73%. Furthermore, the framework achieves a substantial 15.1% reduction in total traveled distance after CMA-ES optimization. The cooperative utilization of UAVs and UGVs effectively balances their complementary advantages, highlighting the system’s scalability and practicality for real-world deployment.
keywords:
@@
1 Introduction
Catastrophic events like earthquakes, floods, and hurricanes frequently destroy critical infrastructure such as roads, bridges, and communication networks, severely hindering traditional emergency response capabilities. In these post-disaster scenarios, providing timely medical assistance becomes an urgent priority, directly impacting survival rates and recovery outcomes. A review of the 2023 Kahramanmaraş earthquakes revealed that significant infrastructure damage impeded rescue efficiency and medical aid, emphasizing the need for better preparedness and more resilient infrastructure[1]. UAVs and UGVs have emerged as promising solutions due to their flexibility, rapid deployment, and ability to navigate hazardous environments inaccessible to human responders. UAVs offer aerial reconnaissance and swift delivery, bypassing ground-level obstacles, while UGVs provide stable transportation for heavier supplies and possess longer operational endurance.
@@ 1 Introduction
Significant attention has been given to multi-robot cooperative systems for disaster relief, as they leverage the complementary strengths of UAVs and UGVs. De Castro et al.[2] investigated UAV-UGV cooperation for coverage tasks in dynamic environments. Krizmancic et al.[3] focused on multi-robot coordination for construction tasks, and Cognetti et al.[4] utilized UAVs for localization to assist UGV navigation. The relevance of such systems in Synthetic Aperture Radar operations has been reviewed by Queralta et al.[5].
@@
For task allocation, various techniques, including Genetic Algorithms (GA) , Particle Swarm Optimization (PSO) , and auction-based mechanisms, have been explored. For example, Li et al.[6] developed an improved GA for multi-robot task assignment , and Dadgar et al.[7] proposed an adaptive PSO for target searching. Elango et al.[8] combined K-means clustering with auction mechanisms for balanced task allocation and López-González et al.[9] applied parallel GAs for formation control.
@@
Similarly, path planning and optimization methods such as RRT*, Probabilistic Roadmap (PRM), and CMA-ES have been widely used to generate collision-free and efficient trajectories. Li et al.[10], for instance, extended RRT* for multi-robot scenarios to account for motion constraints. Other approaches like the PRM method have been combined with optimization techniques, as seen in the hybrid ABC-PRM-EP framework by Kumar and Sikander[11]. Additionally, trajectory optimization methods such as CMA-ES have been introduced to refine initial paths, minimizing travel distance and overall mission completion time. Díez and Aznar[12] compared CMA-ES with other evolution strategies for learning autonomous navigation behaviors in robot swarms. Integrating these task allocation and path planning strategies significantly contributes to the success of cooperative multi-robot disaster relief operations.
Despite these notable advancements, several critical limitations in existing research remain unresolved, presenting key research gaps that this study aims to address.
First, most current cooperative frameworks primarily consider homogeneous teams of unmanned vehicles (either UAVs or UGVs alone). This approach ignores the significant benefits of heterogeneous cooperation and fails to fully integrate and exploit the distinct, complementary strengths of UAVs and UGVs, such as a UAV’s agility and a UGV’s payload capacity.
@@
Second, many existing task allocation methodologies assume static environments and predetermined task sets. This simplification does not reflect the unpredictable nature of real-world disaster scenarios, where new tasks may emerge and the environment is constantly changing. Traditional optimization and heuristic methods often struggle with the high computational complexity of these dynamic, real-time, multi-task scenarios.
@@
Third, trajectory optimization in disaster environments often focuses exclusively on generating feasible paths or minimizing path length, without adequately considering practical constraints like smoothness, continuity, or energy efficiency. Consequently, existing path planning algorithms can produce jagged or inefficient trajectories that are impractical for real-world navigation.
@@
  • We propose an enhanced Genetic Algorithm for dynamic task allocation among a heterogeneous fleet, which is a significant improvement over methods focused on homogeneous teams. The allocation is seamlessly integrated with a trajectory generation process that uses Informed-RRT* for initial pathfinding and CMA-ES for sequence and path optimization. This unified approach ensures that vehicle capabilities are matched to tasks and that the resulting paths are globally coordinated and locally optimized.
  • We introduce a robust methodology that not only generates collision-free paths but also refines them for real-world application. The combination of Informed-RRT* and CMA-ES specifically addresses the limitations of traditional algorithms by optimizing task visiting sequences to reduce total travel distance and mission time. A systematic trajectory post-processing approach is also applied to ensure practical smoothness and continuity, addressing a common shortcoming in existing planning frameworks.
  • Through extensive simulation experiments in realistic post-disaster scenarios, we validate the effectiveness, robustness, and scalability of our proposed framework. The selection of a hybrid evolutionary algorithmic framework (EGA and CMA-ES) over alternatives such as exact solvers (e.g., Gurobi) or deep reinforcement learning (e.g., A2C) is motivated by the need for a balance between solution quality and computational efficiency in dynamic, large-scale disaster scenarios. Exact solvers often suffer from prohibitive computational complexity for real-time applications, while reinforcement learning methods require extensive training and exhibit lower interpretability. We demonstrate significant reductions in total mission completion time and traveled distance compared to traditional strategies like random assignment and K-Means clustering. Critically, our scalability analysis confirms that the proposed approach maintains its efficiency even as the number of tasks and vehicles increases, outperforming baseline methods that show performance degradation under increased complexity.
The remainder of this paper is organized as follows: Section 2 provides a detailed literature review. Section 3 formulates the problem. Section 4 presents our integrated methodology. Section 5 details the simulation results and analysis. Finally, Section 6 concludes the paper and discusses future work.
2 Literature Review
2.1 Multi-Robot Cooperation in Disaster Relief
Multi-robot systems, particularly heterogeneous teams, have emerged as a promising solution for disaster relief operations due to their flexibility, rapid deployment capabilities, and ability to navigate hazardous or inaccessible environments. UAVs excel in aerial reconnaissance and swift delivery, bypassing ground-level obstacles. In contrast, UGVs provide stable transportation for heavier supplies and possess longer operational endurance. The effective cooperation between these platforms leverages their complementary strengths to enhance rescue efficiency and improve survival rates.
Existing studies have explored multi-robot cooperation in various contexts. For example, de Castro et al. investigated UAV-UGV cooperation for coverage tasks in dynamic environments. Krizmancic et al. focused on coordination for automated construction tasks , while Cognetti et al. utilized UAVs for localization to aid UGV navigation. However, a critical limitation in many current frameworks is the focus on homogeneous teams, which fails to fully integrate and exploit the distinct advantages of both UAVs and UGVs. Addressing this gap by developing a framework that can effectively coordinate a heterogeneous fleet is essential for enhancing rescue capabilities.
2.2 Task Allocation Strategies for Multi-Robot Systems
Effective task allocation is a crucial challenge in disaster scenarios, as it directly impacts resource deployment and overall mission efficiency. Various optimization and heuristic techniques have been extensively studied for this purpose, including Genetic Algorithms, Particle Swarm Optimization, and auction-based mechanisms. For instance, Li et al. developed an improved GA for multi-robot task assignment , while López-González et al. applied parallel GAs for formation control. Dadgar et al. proposed an adaptive PSO for target searching , and Elango et al. combined K-means clustering with auction mechanisms for balanced task allocation.
The success of these methods highlights the potential of optimization techniques in solving complex, dynamic problems. Beyond traditional heuristics, advanced machine learning and AI methods have also demonstrated significant potential in modeling and forecasting complex patterns in other domains, which underscores their broad applicability. For instance, neural networks have been successfully used for forecasting thermal coal futures trading volumes, peanut oil price changes, corn cash prices, and house prices[13, 14, 15]. These models often lead to high accuracy and stability, contributing to efficient and low-cost estimations[16, 17, 18, 19, 20]. Similarly, Gaussian process regressions have been effectively applied to predict residential property price indices and steel price indices, often outperforming traditional econometric models[21, 22, 23, 24, 25]. The use of machine learning has also been extended to material science, such as machine learning the magnetocaloric effect in manganites from lattice parameters[26]. While these applications are distinct from task allocation, they demonstrate the power of AI/ML models to solve complex, non-linear problems, a critical characteristic of multi-robot task assignment.
However, a key limitation of many existing task allocation methodologies is the assumption of static environments and predetermined tasks, which is unrealistic for dynamic post-disaster situations. This gap necessitates the development of more robust optimization approaches that can handle the high computational complexity of real-time, dynamic, and multi-task scenarios.
2.3 Trajectory Planning and Optimization Techniques
Path planning is another crucial aspect of multi-robot missions, requiring the generation of collision-free and efficient trajectories in complex, cluttered environments. Algorithms such as RRT*, Probabilistic Roadmap (PRM), and A* search have been widely employed for this purpose. For example, Li et al. extended RRT* for multi-robot scenarios, considering motion constraints. Other hybrid approaches, like the ABC-PRM-EP framework, have combined PRM with optimization techniques to improve efficiency.
Furthermore, trajectory optimization methods are used to refine initial paths and minimize travel distance and mission completion time. For example, CMA-ES has been compared with other evolution strategies for learning autonomous navigation behaviors in robot swarms. The need for advanced optimization is also evident in other fields where complex data patterns require sophisticated modeling. Graph models, for instance, have been used to analyze causal orderings among property and steel product prices, revealing complex dynamics. Similarly, integrated or composite methods, which combine multiple models, have been shown to improve forecasting accuracy for corn cash prices[13, 14, 15].
Despite notable advancements in multi-robot cooperation, task allocation, and trajectory optimization, several critical limitations remain unresolved.
First, most existing cooperative frameworks primarily focus on homogeneous teams of unmanned vehicles, neglecting the substantial benefits derived from a heterogeneous UAV-UGV partnership. The distinct, complementary strengths of UAVs (agility and rapid coverage) and UGVs (enhanced payload capacity and longer endurance) are not yet sufficiently integrated or exploited in current frameworks.
Second, many proposed task allocation methodologies rely on assumptions of static environments and predetermined task sets, which fail to capture the dynamic and unpredictable nature of real-world post-disaster scenarios. Traditional optimization and heuristic methods often struggle with the high computational complexity that arises in these dynamic, multi-task environments.
Third, trajectory optimization for disaster environments frequently prioritizes feasibility or path length minimization without adequately considering practical constraints such as smoothness, continuity, or energy efficiency, which are crucial for actual UAV and UGV operations. As a result, existing planning algorithms often produce jagged or inefficient trajectories that are impractical for real-world deployment.
3 Problem Formulation
Before delving into the specific mathematical formulations, it is crucial to articulate the overarching problem-solving paradigm that guides our approach. The integrated problem of task allocation, trajectory generation, and sequence optimization for a heterogeneous team in a complex environment is inherently high-dimensional and computationally intractable if solved monolithically. To address this challenge, we adopt a structured, hierarchical, and iterative optimization framework, decomposing the complex problem into a series of interdependent sub-problems.
The process begins with a global task allocation using an EGA, which efficiently distributes tasks among the heterogeneous fleet of UAVs and UGVs by considering their distinct capabilities and constraints. This step produces a preliminary assignment that balances the workload. Subsequently, for each vehicle, the Informed-RRT* algorithm is employed to generate an initial, collision-free, and dynamically feasible trajectory that connects its assigned tasks, transforming the abstract assignment into concrete, safe paths. Finally, to refine the efficiency of these paths, the CMA-ES is applied to optimize the visiting sequence of the assigned tasks for each vehicle, minimizing travel distance and time.
This ”Allocate-Plan-Optimize” workflow provides a tractable and effective strategy. The EGA ensures a globally intelligent distribution of work, Informed-RRT* guarantees local feasibility and safety, and CMA-ES performs a fine-grained optimization to eliminate inefficiencies. The outputs of each stage inform and constrain the next, creating a coherent solution pipeline.
@@ 3 Problem Formulation
3.1 Environment Modeling
The operational theatre is conceptualized as a bounded two-dimensional Euclidean space, denoted by $\mathcal{E}\subset\mathbb{R}^{2}$, which represents a disaster-impacted zone, typically an urban or semi-urban landscape significantly damaged by events such as earthquakes or widespread flooding. Within this environment $\mathcal{E}$, several key elements are defined. A primary consideration is the presence of $N$ static obstacles, collected in the set $\mathcal{O}=\{O_{1},O_{2},\dots,O_{N}\}$. These obstacles, geometrically represented (e.g., as polygons or circles), correspond to impassable areas like collapsed structures, extensive debris fields, or flooded regions that impede vehicle movement. The locations and geometries of these major obstacles are assumed to be known or rapidly mappable post-disaster. Distributed across the environment are $M$ discrete task locations, $\mathcal{T}=\{T_{1},T_{2},\dots,T_{M}\}$, each signifying a point where a specific medical rescue action is required, such as the delivery of critical supplies or site assessment. For the scope of this work, these task locations are assumed to remain fixed throughout the planning phase. Finally, a central base station $B$ is situated within $\mathcal{E}$, serving as the designated origin and final destination for all unmanned vehicles involved in the operation, functioning as the mission control and logistics hub. Consequently, the navigable area for the vehicles, or free space, is $\mathcal{E}_{free}=\mathcal{E}\setminus\bigcup_{i=1}^{N}O_{i}$. A fundamental requirement is that all vehicle movement must be confined to $\mathcal{E}_{free}$, ensuring mandatory collision avoidance with respect to all obstacles in $O_{i}\in\mathcal{O}$. While real-world disaster zones are inherently dynamic, this study adopts a static representation of major obstacles and task points for planning tractability during a given mission cycle. The schematic diagram is shown in Fig. 1.
Figure 1: Schematic diagram of the UAV-UGV collaborative mission environment
3.2 UAV and UGV Operational Constraints
@@
where $P_{i}$ denotes the waypoints along the planned trajectory generated by the path planning algorithm (e.g., Informed-RRT*), $d(\cdot,\cdot)$ is the Euclidean distance, and $D_{\text{UAV,max}}$ is the maximum flight distance.
@@
While this constraint is formulated using the sum of Euclidean distances for simplicity and computational tractability, it is acknowledged that the actual energy consumption of UAVs and UGVs is more accurately represented by the total length of the final, dynamically feasible trajectory, which incorporates the effects of path curvature, kinematic constraints, and environmental factors. The total trajectory length serves as a better proxy for real-world energy expenditure than the sum of straight-line segments. In future work, we aim to integrate more sophisticated energy models that account for vehicle dynamics and terrain characteristics to further refine the planning process.
@@
3.3 Task Definition and Constraints
@@ 3.4 Optimization Objectives
3.5 Problem Scope and Formal Summary
In essence, the problem tackled in this paper is the optimization of task allocation and trajectory planning for a heterogeneous fleet $\mathcal{V}$ of UAVs and UGVs operating in a post-disaster setting. The operational context is defined by an environment $\mathcal{E}$ containing a set of known, static obstacles $\mathcal{O}$, a designated set of medical rescue task locations $\mathcal{T}$, and a central base station $B$. The core objective is to determine an optimal operational plan, which encompasses both the assignment of each task $T_{j}\in\mathcal{T}$ exclusively to a single vehicle $V_{i}\in\mathcal{V}$ (represented by assignment $X$), and the subsequent generation of an ordered task sequence $\pi_{i}$ and a feasible, collision-free trajectory $\mathcal{P}_{i}(t)$ for each vehicle. Each trajectory must originate from $B$, visit the assigned tasks in the determined sequence $\pi_{i}$, and terminate back at $B$, while strictly adhering to all vehicle-specific constraints, particularly maximum travel endurance and payload limits. The quality of the plan is primarily evaluated based on minimizing the overall mission completion time and the total distance traveled by the entire fleet. Critically, the return leg from the final task location back to the base station $B$ is planned as a distinct, independent path segment using the Informed-RRT* algorithm, ensuring an optimal and collision-free route, rather than simply reversing the outbound path.